I need help with this proportion I have 9 other problems like this
7/9=b/b-10

Answer by Alan3354(69443)

I need help with this proportion I have 9 other problems like this
If you mean:
7/9=b/(b-10)
Cross multiply
7*(b-10) = 9b
7b-70 = 9b
-70 = 2b
b = -35